### v1

# CogDDN: A Cognitive Demand-Driven Navigation with Decision Optimization and Dual-Process Thinking

Abstract.
Mobile robots are increasingly required to navigate and interact within unknown and unstructured environments to meet human demands. Demand-driven navigation (DDN) enables robots to identify and locate objects based on implicit human intent, even when object locations are unknown. However, traditional data-driven DDN methods rely on pre-collected data for model training and decision-making, limiting their generalization capability in unseen scenarios. In this paper, we propose CogDDN, a VLM-based framework that emulates the human cognitive and learning mechanisms by integrating fast and slow thinking systems and selectively identifying key objects essential to fulfilling user demands. CogDDN identifies appropriate target objects by semantically aligning detected objects with the given instructions. Furthermore, it incorporates a dual-process decision-making module, comprising a Heuristic Process for rapid, efficient decisions and an Analytic Process that analyzes past errors, accumulates them in a knowledge base, and continuously improves performance. Chain of Thought (CoT) reasoning strengthens the decision-making process. Extensive closed-loop evaluations on the AI2Thor simulator with the ProcThor dataset show that CogDDN outperforms single-view camera-only methods by 15%, demonstrating significant improvements in navigation accuracy and adaptability. The project page is available at https://yuehaohuang.github.io/CogDDN/.
1. Introduction
Mobile robots are increasingly essential for assisting humans in unknown and unstructured environments, such as households, hospitals, and warehouses. To operate effectively, they must be capable of interpreting and responding to implicit human demands, even when target objects are uncertain or not explicitly specified. For instance, a hungry person instinctively seeks food based on internal cues rather than direct instructions. Similarly, demand-driven navigation (DDN) (Wang et al., 2023, 2024c) enables robots to locate objects that fulfill such implicit needs without relying on predefined object lists or locations. However, as shown in Figure 1 (a), conventional data-driven DDN methods (Wang et al., 2023, 2024c) depend heavily on extensive pre-collected data for model training and decision-making, limiting their generalization to novel scenarios and vague instructions.
Comparison between cognition-driven and data-driven.
Recent breakthroughs in Large Language Models (LLMs) and Vision-Language Models (VLMs) have demonstrated strong reasoning capabilities and extensive world knowledge, making them well-suited for navigation tasks. Cognition-driven approaches (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025; Wang et al., 2024b; Shah et al., 2022) leverage these models to help agents interpret instructions and locate target objects. The DDN tasks combine natural language with visual perception, requiring agents to infer and act upon implicit user intent. As illustrated in Figure 1 (b), our framework integrates LLMs and VLMs for DDN, enabling more sophisticated reasoning over vague or ambiguous instructions. Additionally, VLMs improve navigation by jointly processing visual and linguistic inputs, effectively bridging semantic intent with spatial context. This cognition-driven, multi-agent approach enhances the agent’s spatial and contextual understanding, resulting in more accurate and adaptable navigation. Figure 1 (c) shows that the cognition-driven method generalizes better than traditional data-driven approaches, exhibiting a smaller performance drop between seen and unseen scenarios.
CogDDN Framework for Demand-Driven Navigation Tasks.
This paper presents CogDDN, a novel VLM-based approach inspired by the dual-process framework for closed-loop decision-making. A key distinction is that it emulates human cognitive and learning mechanisms. Unlike many VLM-based navigation systems (Long et al., 2024; Cai et al., 2024) that rely on additional views, depth, or maps, our system uses only front-facing views, aligning with human search behavior and real-world robotic constraints. At a result, our system performs a more realistic navigation task by mirroring the perceptual constraints and decision-making processes of human object search.
In parallel with human decision-making, CogDDN incorporates a demand-matching module that leverages LLMs to accurately link instructions to relevant objects, allowing the agent to identify and locate key items that satisfy implicit user needs. Furthermore, we introduce a dual-process decision-making system based on dual-process theory (Kahneman, 2011; Evans and Stanovich, 2013; Wason and Evans, 1974), which simulates human cognition through two components: a Heuristic Process (System-I) and an Analytic Process (System-II) (Mei et al., 2024; Ma et al., 2025). The Heuristic Process enables rapid, intuitive decisions based on prior knowledge, while the Analytic Process refines these decisions by analyzing new data and updating the knowledge base. Before navigation, we construct a knowledge base to store high-quality strategies adaptable to diverse scenarios. The Heuristic Process utilizes this accumulated knowledge, refined through supervised fine-tuning (SFT), to guide initial decisions. Meanwhile, the Analytic Process updates the knowledge base when encountering obstacles or challenges, enabling continuous learning and improved decision-making. In contrast to direct end-to-end methods, our system integrates Chain of Thought (CoT) (Wei et al., 2022) reasoning, enhancing the agent’s ability to reason through complex scenarios and make informed decisions, rather than merely outputting final actions.
The main contributions are summarized as follows:
We propose CogDDN, a novel VLM-based framework for DDN tasks that simulates human-like cognitive mechanisms and learning processes.
We introduce a dual-process decision-making module. Without human involvement, our approach allows the empirical Heuristic Process to inherit the capabilities of the rational Analytic Process through a self-supervised mechanism.
CogDDN uses the Analytic Process and a knowledge accumulation mechanism to enrich a transferable knowledge base, enabling continual learning and effective generalization to novel navigation environments.
2. Related Works
2.1. Demand-driven Navigation
Demand-driven Navigation (DDN) (Wang et al., 2023, 2024c) is an innovative task requiring an agent to navigate an environment based on high-level human demands described in natural language. Unlike traditional visual navigation tasks like Visual Object Navigation (VON) (Ramakrishnan et al., 2022; Zhao et al., 2023; Kiran et al., 2022; Chaplot et al., 2020; Gadre et al., 2022; Pal et al., 2021) and Visual Language Navigation (VLN) (Wu et al., 2024; Yue et al., 2024; Zhang et al., 2024b; Zhou et al., 2024a, b; Hong et al., 2025; Goetting et al., 2024; Cai et al., 2024), DDN does not focus on specific object categories or step-by-step instructions but rather on fulfilling user-defined needs. Using abstract, high-level instructions, DDN enables the agent to infer which objects in the environment are relevant to the given demand. This makes DDN more flexible and dynamic than VON or VLN, as it centers on addressing ”human demands” rather than specific objects. The agent must interpret these demands and make inferences from the current scene, providing a more generalized approach to navigation. The main challenge of DDN lies in interpreting and fulfilling human demand within the visual context.
2.2. Transitioning from Data-Driven to Cognition-Driven Navigation
Early navigation methods were predominantly data-driven, relying on imitation learning (Nguyen et al., 2019; Kamath et al., 2023) and reinforcement learning (Wang et al., 2018; Chen et al., 2022) to train reactive policies on large-scale labeled datasets. While these approaches achieved promising results within their training distributions, they often failed to generalize to out-of-distribution scenarios (Cai et al., 2025; Zhan et al., 2024), mainly due to limited semantic reasoning and contextual awareness, and often incur high computational costs during both training and deployment (Zhang et al., 2024c; Wei et al., 2024; Xiang and Zhang, 2024).
Motivated by the success of pre-trained vision-language models in various multimodal reasoning tasks (Ma et al., 2024; Wang et al., 2024d; Lu et al., 2025; Liu et al., 2025a, b), recent navigation research has begun to adopt these models as cognitive priors to move toward more robust and human-like navigation strategies (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025). Inspired by human-like reasoning, these systems unify perception and language to support language-driven goal specification (Mirjalili et al., 2024) and semantic understanding (Bangguo Yu and Cao, 2024). Unlike traditional methods that depend on structured instructions or spatial maps, VLM-based approaches allow agents to interpret open-ended natural language commands and navigate in complex, unfamiliar environments (Zhang et al., 2024a; Song et al., 2025). These systems emphasize multimodal grounding, context-aware planning, and continual learning, and they often incorporate reasoning techniques such as Chain-of-Thought to support interpretable and generalizable behavior.
A detailed architecture diagram of the proposed CogDDN.
3. Method
In this section, we describe the design of CogDDN, a closed-loop, anthropomorphic indoor navigation system. As shown in Figure 3, the system consists of five key components: 3D Robot Perception for identifying objects, an LLM for Demand Matching (Section 3.2), a Knowledge Base for storing experiences (Section 3.3), and a dual-process decision-making module with the Heuristic Process (Section 3.4) and the Analytic Process (Section 3.5). We begin by formally defining the DDN task in Section 3.1, followed by detailed descriptions of each component.
3.1. Task Definition
CogDDN tasks require agents to locate relevant objects in a given environment based on human demands and perceptual inputs. They can be formalized as a Partially Observable Markov Decision Process (POMDP), defined as
where represents the state space (the agent’s current state), is the observation space (including human demands and images), is the action space (e.g., MoveAhead, RotateLeft, LookDown, etc.), and is the action execution function. For instance, a user may state, ”I am thirsty.” In response, the agent must identify an object capable of quenching thirst based on the current visual observation and then execute the actions to locate the object. After each action, the agent receives updated observations and generates new actions based on the updated information to continue searching for the object. The task is complete once the agent is within 1.5 meters of the target object.
3.2. Demand Matching
Objects that fulfill the same human demand in demand-driven navigation share key properties. For instance, paintings, houseplants, and statues are suitable candidates for decorating a space, as they enhance visual appeal and complement the intended environment. This relationship between human demands and object properties is grounded in universal knowledge.
While LLMs are proficient in reasoning based on instructions and properties, they encounter difficulties when an exact match for a request is unavailable. In such cases, LLMs may suggest suboptimal objects. For instance, if a user requests, ”I need something to hold my flowers,” a mug might be suggested despite being a less-than-ideal choice. This tendency significantly reduces the model’s accuracy in selecting the most suitable object. To mitigate this issue, we apply supervised fine-tuning (SFT) to train the LLM, allowing it better to align objects with user requirements. This method improves the model’s capacity to identify and recommend the most appropriate objects, ensuring it avoids suggesting less optimal options, even when an exact match is not readily apparent.
Formally, the Demand Matching module implements a target objects generation function , where represents the space of observation space including human demands and images, is the space of demand properties, and is the space of target objects. By leveraging a fine-tuned LLM for demand matching, the system benefits from its ability to reason over complex instructions and object properties. This capability prevents it from forcing matches between human demands and objects when an exact match is unavailable, ensuring more appropriate suggestions even in ambiguous scenarios. Please refer to Appendix A.2 for annotation details.
The Construction of the Knowledge Base.
3.3. Knowledge Base
In indoor navigation scenarios, humans typically prioritize the approximate location of the target object and the scene description, focusing on navigable areas. This approach helps mitigate information overload, enhances reaction times, and reduces cognitive load. Consequently, we use scene descriptions to delineate passable areas, providing a sufficient contextual understanding of the environment. Additionally, Chain of Thought (CoT) reasoning enhances the model’s inference process by leveraging the visual model’s language-based reasoning capabilities. This approach contrasts with direct end-to-end methods by explicitly incorporating reasoning steps before generating final action decisions.
Inspired by LeapVAD (Ma et al., 2025), we generate additional data with outputs divided into three components: 1) Scene Description: This component outlines the environment based on the current viewpoint image, highlighting passable areas and task-related objects that may affect mobility. 2) Reasoning: This component leverages the visual-language model’s spatial awareness and commonsense reasoning capabilities based on the target’s location and the scene description to guide the agent’s navigation. 3) Decision: Drawing from the output of the reasoning component, the model generates the optimal decision for the current navigation task, ensuring the selection of the most effective action. This process incorporates knowledge transfer mechanisms to enhance navigation performance.
Building on this foundational approach of leveraging scene descriptions and reasoning for navigation optimization, we construct a knowledge base that integrates environmental data with dynamic decision-making processes. As illustrated in Figure 4, during the execution of the trajectory generated by the A* algorithm (Hart et al., 1968), once the target object is detected, the current viewpoint image , target object , and final action S are input into the VLM to generate scene descriptions D and reasoning R, which are then recorded as an experience in the knowledge base. This accumulated experience can be progressively transferred to the Heuristic Process. As detailed in Section 3.4, we utilize the collected data {} to train the Heuristic Process’s ability to follow instructions and adhere to formatting standards. Please refer to Figure 9 in Appendix A.3 for the data format.
Detailed procedure of the heuristic process.
3.4. Heuristic Process
While the VLM effectively generates reasoning based on decisions from the A* algorithm (Hart et al., 1968), it struggles to directly infer the relationship between reasoning and decisions. Moreover, its relatively slow processing speed often leads to redundancy and repetition, limiting practical use in indoor navigation. Formally, the Heuristic Process module implements an action execution function , where represents the space of target objects, depicts the state space, and is the action space, which is presented in table 1. Inspired by human navigation, where efficiency is gained through repeated practice and environmental familiarity, allowing individuals to navigate with minimal cognitive effort, we introduce the Heuristic Process in CogDDN. As demonstrated in Figure 5, it consists of two modules: Explore and Exploit, both leveraging Chain-of-Thought (CoT) reasoning to adapt to new environments and optimize learned navigation execution. The Heuristic Process is expressed as:
Explore. The Explore module is activated when the Demand Matching module cannot identify a suitable target within the current scene. In such cases, the system searches the environment by generating exploratory actions based on its visual perspective, utilizing the VLM. This aims to uncover previously overlooked objects or paths. As Figure 2 shows, the system first generates a Scene Description D based on the current viewpoint image . Subsequently, it conducts Reasoning R to direct the exploration process. The system’s Decision S consists of a series of actions to explore unvisited regions. The module incorporates information from prior actions and recent rotations to minimize redundant exploration. If obstacles hinder progress, the flag is set to true, prompting the system to dynamically generate new exploratory actions, ensuring continuous progress without unnecessary repetition or resource depletion.
| Action | Definition |
| MoveAhead | Move forward by meters. |
| RotateLeft | Rotate the agent to the left. |
| RotateRight | Rotate the agent to the right. |
| LookUp | Tilt the agent’s camera upward by . |
| LookDown | Tilt the agent’s camera downward by . |
| Done | Indicate that the goal has been found. |
Exploit. Once the Demand Matching module identifies a target object, the system transitions to the Exploit module. Unlike the Explore phase, which focuses on broader navigation and search strategies, the Exploit phase emphasizes precise, deliberate actions to achieve the navigation goal. This phase leverages a fine-tuned version of the VLM, optimized through prior exploration and experiences stored in the knowledge base. Fine-tuning improves the model’s ability to make accurate predictions and decisions based on subtle environmental cues, enabling the system to handle complex tasks with high precision. Guided by CoT, the system first generates a Scene Description D based on the current viewpoint image and the information of target objects , performs Reasoning R to infer the most suitable course of action based on the D and , and produces the most optimal Decision S, a single action focused on reaching the target. Focusing on one action rather than multiple potential actions improves efficiency and computational performance, ensuring rapid and effective goal achievement.
Detailed procedure of the reflection mechanism.
| Method | Seen Scene | Unseen Scene | ||||||||||
| Seen Ins. | Unseen Ins. | Seen Ins. | Unseen Ins. | |||||||||
| NSR | SPL | SSR | NSR | SPL | SSR | NSR | SPL | SSR | NSR | SPL | SSR | |
| Random (Wang et al., 2023) | 5.2 | 2.6 | 3.0 | 3.7 | 2.6 | 2.3 | 4.8 | 3.3 | 2.8 | 3.5 | 1.9 | 1.4 |
| VTN-demand (Wang et al., 2023) | 6.3 | 4.2 | 3.2 | 5.2 | 3.1 | 2.8 | 5.0 | 3.2 | 2.8 | 6.6 | 4.0 | 3.3 |
| VTN-CLIP-demand (Wang et al., 2023) | 12.0 | 5.1 | 5.7 | 10.7 | 3.5 | 5.0 | 10.0 | 3.6 | 4.0 | 9.3 | 3.9 | 4.0 |
| VTN-GPT* (Wang et al., 2023) | 1.6 | 0.5 | 0 | 1.4 | 0.4 | 0.5 | 1.3 | 0.2 | 0.3 | 0.9 | 0.4 | 0.5 |
| ZSON-demand (Wang et al., 2023) | 4.2 | 2.7 | 1.9 | 4.6 | 3.1 | 2.0 | 4.1 | 2.9 | 1.2 | 3.5 | 2.4 | 1.1 |
| ZSON-GPT (Wang et al., 2023) | 4.0 | 1.1 | 0.3 | 3.6 | 1.9 | 0.3 | 2.5 | 0.7 | 0.2 | 3.2 | 0.9 | 0.2 |
| CLIP-Nav-MiniGPT-4 (Wang et al., 2023) | 4.0 | 4.0 | 2.0 | 3.0 | 3.0 | 2.0 | 4.0 | 3.7 | 2.0 | 5.0 | 5.0 | 3.0 |
| CLIP-Nav-GPT* (Wang et al., 2023) | 5.0 | 5.0 | 4.0 | 6.0 | 5.5 | 5.0 | 5.5 | 5.3 | 4.0 | 4.0 | 3.0 | 2.0 |
| FBE-MiniGPT-4 (Wang et al., 2023) | 3.5 | 3.0 | 2.2 | 3.5 | 3.2 | 2.0 | 3.5 | 3.5 | 2.0 | 4.0 | 4.0 | 3.5 |
| FBE-GPT* (Wang et al., 2023) | 5 | 4.3 | 4.3 | 5.5 | 5.0 | 5.5 | 4.5 | 4.3 | 4.5 | 5.5 | 5.0 | 5.5 |
| GPT-3-Prompt* (Wang et al., 2023) | 0.3 | 0.01 | 0 | 0.3 | 0.01 | 0 | 0.3 | 0.01 | 0 | 0.3 | 0.01 | 0 |
| MiniGPT-4 (Wang et al., 2023) | 2.9 | 2.0 | 2.5 | 2.9 | 2.0 | 2.5 | 2.9 | 2.0 | 2.5 | 2.9 | 2.0 | 2.5 |
| DDN (Wang et al., 2023) | 21.5 | 9.8 | 7.5 | 19.3 | 9.4 | 4.5 | 14.2 | 6.4 | 5.7 | 16.1 | 8.4 | 6.0 |
| CogDDN(ours) | 38.3 | 17.2 | 29.8 | 37.5 | 18.0 | 28.6 | 33.3 | 16.4 | 27.1 | 34.5 | 17.1 | 27.5 |
3.5. Analytic Process
The Analytic Process is employed to reflect on navigation obstacles, as illustrated in Figure 6. Through comprehensive pre-training on diverse datasets, VLMs naturally accumulate broad world knowledge, enabling them to address complex challenges with sophisticated reasoning and insight. This capability aligns with the objectives of the Analytic Process, which relies on detailed analysis and contextual understanding to draw sound inferences in indoor navigation.
Specifically, in a closed-loop indoor navigation scenario where the Demand Matching and Heuristic Process modules operate, any obstruction encountered during navigation will trigger the reflection mechanism. The Analytic Process utilizes the vast world knowledge embedded within the VLMs to analyze the situation based on previously gathered information. During this process, the object location, scene description, reasoning, and decision from the preceding frame before the obstruction are forwarded to the Analytic Process. The system then carefully analyzes the cause of the event, identifies the error, and generates corrected reasoning and decisions. The insights gained from this reflection process are subsequently integrated into the knowledge base, enabling the system to learn from failures continuously. This iterative learning process progressively enhances the system’s decision-making capabilities, resulting in more informed and accurate navigation strategies for future tasks.
4. Experiments
4.1. Data preparation
The experiments use the AI2Thor (Kolve et al., 2017) and the ProcThor dataset (Deitke et al., 2022). Following DDN (Wang et al., 2023), we evaluate 600 scenes (200 each from ProcThor’s train, validation, and test sets). These environments contain 109 distinct object categories, each capable of fulfilling the demand instructions. For instructions, we use the 1692 training, 241 validation, and 485 test instructions provided by DDN (Wang et al., 2023).
Data for 3D Robot Perception. We used data from the AI2Thor simulator to train the 3D Robot Perception Module. Specifically, we collected 55K frames, ensuring that 70% of the data was allocated for training, 10% for validation, and 20% for testing. For clarity, we provide details and methods about data in Appendix A.1.
Data for Demand Matching. By leveraging the instruction data, we curated an instruction-attribute dataset for the Demand Matching Module. The dataset was constructed by inputting instructions and associated objects into the LLM, which generated common attributes shared by objects fulfilling the instructions. We collected 7.2K data for training, 1.4K for validation, and 2.1K for testing. Additional details are provided in Appendix A.2.
Data for Heuristic Process. To support the supervised fine-tuning of our Heuristic Process using the knowledge base, we curated datasets by collecting data from the AI2Thor simulator. The construction of the knowledge base is discussed in Section 3.3. We generated approximately 72K trajectories, with 80% used for training and 20% reserved for validation. Additional details are provided in Appendix A.3.
4.2. Implementation Details
We employ UniMODE (Li et al., 2024) for 3D robot perception and GPT-4 for the Analytic Process to enable rational and logical thinking. Qwen2-VL-7B (Wang et al., 2024a) is employed for both the Exploit module in the Heuristic Process and Demand Matching, with multiple LoRA configurations for optimal performance. All experiments are trained on 4 RTX6000 GPUs. Please refer to Appendix D for further implementation details.
4.3. Evaluation in Closed-Loop Navigation
We conduct closed-loop experiments in AI2Thor, a widely-used open-source simulator, to evaluate the performance of CogDDN on 400 scenes using three metrics: Navigation Success Rate (NSR), Navigation Success Rate weighted by the path length (SPL) (Anderson et al., 2018), and Selection Success Rate (SSR). Table 2 compares our method with competitive methods offered by DDN (Wang et al., 2023) on the ProcThor benchmark.
As shown in Table 2, CogDDN outperforms all single-view camera-only methods. With the LLM’s reasoning capabilities, the Demand Matching module exhibits high generalization, with minimal performance gaps between seen and unseen instructions. The main gap between seen and unseen scenes stems from the 3D Robot Perception. Additionally, as shown in Figure 1 (c), CogDDN demonstrates a smaller relative drop between seen and unseen conditions, proving its strong generalization ability. Furthermore, as shown in Table 3, we compare CogDDN with SOTA methods InstructNav (Long et al., 2024) on unseen scenes and instructions. Remarkably, CogDDN achieves comparable performance, even though InstructNav leverages depth maps as additional input.
4.4. Ablation Study
We conduct ablation studies on Exploit, CoT, and Reflection in a closed-loop navigation system. The experiments were carried out on both unseen scenes and unseen instructions, demonstrating the generalization and continuous learning capabilities of our CogDDN.
Ablation Study of Exploit. The Exploit module uses a fine-tuned VLM to generate precise, deliberate single-step actions critical for approaching the target. To assess the impact of fine-tuning, we replace the Exploit module with a vanilla GPT-4 model not adapted for navigation and evaluate performance in two settings: (i) generating one action per step and (ii) generating a full action sequence. As shown in Table 4, this substitution causes a significant performance drop, underscoring the importance of task-specific fine-tuning for accurate, step-wise decisions in the Exploit phase.
| Method | Modality | Reflection | NSR | SPL |
| InstructNav (Long et al., 2024) | D+I | 30.0 | 14.2 | |
| DDN (Wang et al., 2023) | I | 16.1 | 8.4 | |
| CogDDN(ours) | I | 34.5 | 17.1 |
| Method | Single Action | NSR | SPL | SSR |
| CogDDN(w/o Exploit) | 24.2 | 13.9 | 19.2 | |
| CogDDN(w/o Exploit) | 29.1 | 12.4 | 20.2 | |
| CogDDN(w/o CoT) | 27.3 | 11.5 | 21.1 | |
| CogDDN | 34.5 | 17.1 | 27.5 |
Ablation study of CoT. In both the Heuristic and Analytic Processes, we use CoT to guide the VLM in generating high-quality decisions. To evaluate its effect, we remove CoT and generate decisions directly. As shown in Table 4, this results in a significant performance drop, underscoring its importance within the dual-process framework. In the Heuristic Process, CoT helps leverage environmental information for decision-making. In the Analytic Process, CoT enables the VLM to analyze obstacles and uncover root causes, resulting in more accurate and informed decisions.
Ablation study of Reflection. Reflection plays a crucial role in enhancing the continuous learning capabilities of our proposed CogDDN. It involves analyzing mistakes and integrating correct experiences into the knowledge base, thereby promoting self-improvement through proactive summarization and the accumulation of experiences in previously unseen environments. In our closed-loop experiment, we conducted four rounds of 500 epochs each, where the experiences generated by the reflection mechanism were added to the knowledge base and used to fine-tune the VLM. As shown in Figure 7, after four rounds of iterations, both SSR and NSR showed marginal improvements, as the reflection mechanism allows the agent to navigate around obstacles once they are encountered, resulting in only modest improvements in these metrics. However, SPL exhibits a significant increase, as the agent, learning from experiences, anticipates obstacles and takes proactive actions. This highlights the continuous learning capability of our CogDDN, where performance improves with experience.
Ablation of Reflection.
5. Conclusion
In this paper, we present CogDDN, a dual-process closed-loop navigation system that mimics the human attentional mechanism by selectively focusing on key objects necessary for fulfilling user demands. We have developed a dual-process decision-making module that simulates human cognitive processes. Chain of Thought (CoT) reasoning is incorporated to enhance the decision-making process. Moreover, we have integrated a cumulative knowledge base to enable the continuous self-improvement of the algorithm. Our evaluations demonstrate that the proposed method achieves remarkable closed-loop performance while requiring limited training data.
Limitations and Broader Impacts. The current CogDDN system incorporates only short-term memory in the Explore module, lacking long-term memory to identify previously explored areas. Additionally, employing GPT-4 as a decision-making module in the Explore phase is computationally expensive and impractical for real-world deployment. Moreover, updating the Heuristic Process requires SFT whenever new experiences are added to the Knowledge Base, leading to inefficiency. In the future, we will design a new system focusing on long-term memory and end-to-end navigation.
Acknowledgements.
This work is supported by the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2024A09).References
- (1)
- Anderson et al. (2018) Peter Anderson, Angel Chang, Devendra Singh Chaplot, Alexey Dosovitskiy, Saurabh Gupta, Vladlen Koltun, Jana Kosecka, Jitendra Malik, Roozbeh Mottaghi, Manolis Savva, et al. 2018. On evaluation of embodied navigation agents. arXiv preprint arXiv:1807.06757 (2018).
- Bangguo Yu and Cao (2024) Lei Han Hamidreza Kasaei Tingguang Li Bangguo Yu, Yuzhen Liu and Ming Cao. 2024. VLN-Game: Vision-Language Equilibrium Search for Zero-Shot Semantic Navigation. arXiv:2411.11609 (2024).
- Cai et al. (2024) Wenzhe Cai, Siyuan Huang, Guangran Cheng, Yuxing Long, Peng Gao, Changyin Sun, and Hao Dong. 2024. Bridging zero-shot object navigation and foundation models through pixel-guided navigation skill. In 2024 IEEE International Conference on Robotics and Automation (ICRA). IEEE, 5228–5234.
- Cai et al. (2025) Yuxin Cai, Xiangkun He, Maonan Wang, Hongliang Guo, Wei-Yun Yau, and Chen Lv. 2025. CL-CoTNav: Closed-Loop Hierarchical Chain-of-Thought for Zero-Shot Object-Goal Navigation with Vision-Language Models. arXiv preprint arXiv:2504.09000 (2025).
- Chaplot et al. (2020) Devendra Singh Chaplot, Dhiraj Prakashchand Gandhi, Abhinav Gupta, and Russ R Salakhutdinov. 2020. Object goal navigation using goal-oriented semantic exploration. Advances in Neural Information Processing Systems 33 (2020), 4247–4258.
- Chen et al. (2022) Jinyu Chen, Chen Gao, Erli Meng, Qiong Zhang, and Si Liu. 2022. Reinforced structured state-evolution for vision-language navigation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 15450–15459.
- Deitke et al. (2022) Matt Deitke, Eli VanderBilt, Alvaro Herrasti, Luca Weihs, Kiana Ehsani, Jordi Salvador, Winson Han, Eric Kolve, Aniruddha Kembhavi, and Roozbeh Mottaghi. 2022. ProcTHOR: Large-Scale Embodied AI Using Procedural Generation. Advances in Neural Information Processing Systems 35 (2022), 5982–5994.
- Evans and Stanovich (2013) Jonathan St BT Evans and Keith E Stanovich. 2013. Dual-process theories of higher cognition: Advancing the debate. Perspectives on psychological science 8, 3 (2013), 223–241.
- Gadre et al. (2022) Samir Yitzhak Gadre, Mitchell Wortsman, Gabriel Ilharco, Ludwig Schmidt, and Shuran Song. 2022. Clip on wheels: Zero-shot object navigation as object localization and exploration. arXiv preprint arXiv:2203.10421 3, 4 (2022), 7.
- Goetting et al. (2024) Dylan Goetting, Himanshu Gaurav Singh, and Antonio Loquercio. 2024. End-to-End Navigation with Vision Language Models: Transforming Spatial Reasoning into Question-Answering. arXiv preprint arXiv:2411.05755 (2024).
- Hart et al. (1968) Peter E Hart, Nils J Nilsson, and Bertram Raphael. 1968. A formal basis for the heuristic determination of minimum cost paths. IEEE transactions on Systems Science and Cybernetics 4, 2 (1968), 100–107.
- Hong et al. (2025) Haodong Hong, Yanyuan Qiao, Sen Wang, Jiajun Liu, and Qi Wu. 2025. General Scene Adaptation for Vision-and-Language Navigation. arXiv preprint arXiv:2501.17403 (2025).
- Kahneman (2011) Daniel Kahneman. 2011. Fast and slow thinking. Allen Lane and Penguin Books, New York (2011).
- Kamath et al. (2023) Aishwarya Kamath, Peter Anderson, Su Wang, Jing Yu Koh, Alexander Ku, Austin Waters, Yinfei Yang, Jason Baldridge, and Zarana Parekh. 2023. A new path: Scaling vision-and-language navigation with synthetic instructions and imitation learning. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 10813–10823.
- Kiran et al. (2022) DA Sasi Kiran, Kritika Anand, Chaitanya Kharyal, Gulshan Kumar, Nandiraju Gireesh, Snehasis Banerjee, Ruddra Dev Roychoudhury, Mohan Sridharan, Brojeshwar Bhowmick, and Madhava Krishna. 2022. Spatial relation graph and graph convolutional network for object goal navigation. In 2022 IEEE 18th International Conference on Automation Science and Engineering (CASE). IEEE, 1392–1398.
- Kolve et al. (2017) Eric Kolve, Roozbeh Mottaghi, Winson Han, Eli VanderBilt, Luca Weihs, Alvaro Herrasti, Matt Deitke, Kiana Ehsani, Daniel Gordon, Yuke Zhu, et al. 2017. Ai2-thor: An interactive 3d environment for visual ai. arXiv preprint arXiv:1712.05474 (2017).
- Li et al. (2024) Zhuoling Li, Xiaogang Xu, SerNam Lim, and Hengshuang Zhao. 2024. UniMODE: Unified monocular 3D object detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 16561–16570.
- Liu et al. (2025a) Guangyi Liu, Pengxiang Zhao, Liang Liu, Zhiming Chen, Yuxiang Chai, Shuai Ren, Hao Wang, Shibo He, and Wenchao Meng. 2025a. Learnact: Few-shot mobile gui agent with a unified demonstration benchmark. arXiv preprint arXiv:2504.13805 (2025).
- Liu et al. (2025b) Guangyi Liu, Pengxiang Zhao, Liang Liu, Yaxuan Guo, Han Xiao, Weifeng Lin, Yuxiang Chai, Yue Han, Shuai Ren, Hao Wang, et al. 2025b. Llm-powered gui agents in phone automation: Surveying progress and prospects. arXiv preprint arXiv:2504.19838 (2025).
- Long et al. (2024) Yuxing Long, Wenzhe Cai, Hongcheng Wang, Guanqi Zhan, and Hao Dong. 2024. Instructnav: Zero-shot system for generic instruction navigation in unexplored environment. arXiv preprint arXiv:2406.04882 (2024).
- Lu et al. (2025) Zhengxi Lu, Yuxiang Chai, Yaxuan Guo, Xi Yin, Liang Liu, Hao Wang, Han Xiao, Shuai Ren, Guanjing Xiong, and Hongsheng Li. 2025. UI-R1: Enhancing Efficient Action Prediction of GUI Agents by Reinforcement Learning. arXiv preprint arXiv:2503.21620 (2025).
- Ma et al. (2024) Yukai Ma, Jianbiao Mei, Xuemeng Yang, Licheng Wen, Weihua Xu, Jiangning Zhang, Xingxing Zuo, Botian Shi, and Yong Liu. 2024. Licrocc: Teach radar for accurate semantic occupancy prediction using lidar and camera. IEEE Robotics and Automation Letters (2024).
- Ma et al. (2025) Yukai Ma, Tiantian Wei, Naiting Zhong, Jianbiao Mei, Tao Hu, Licheng Wen, Xuemeng Yang, Botian Shi, and Yong Liu. 2025. LeapVAD: A Leap in Autonomous Driving via Cognitive Perception and Dual-Process Thinking. arXiv preprint arXiv:2501.08168 (2025).
- Mei et al. (2024) Jianbiao Mei, Yukai Ma, Xuemeng Yang, Licheng Wen, Xinyu Cai, Xin Li, Daocheng Fu, Bo Zhang, Pinlong Cai, Min Dou, et al. 2024. Continuously learning, adapting, and improving: A dual-process approach to autonomous driving. arXiv preprint arXiv:2405.15324 (2024).
- Mirjalili et al. (2024) Reihaneh Mirjalili, Michael Krawez, Florian Walter, and Wolfram Burgard. 2024. VLM-Vac: Enhancing Smart Vacuums through VLM Knowledge Distillation and Language-Guided Experience Replay. ArXiv abs/2409.14096 (2024). https://api.semanticscholar.org/CorpusID:272826708
- Nguyen et al. (2019) Khanh Nguyen, Debadeepta Dey, Chris Brockett, and Bill Dolan. 2019. Vision-based navigation with language-based assistance via imitation learning with indirect intervention. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 12527–12537.
- Nie et al. (2025) Dujun Nie, Xianda Guo, Yiqun Duan, Ruijun Zhang, and Long Chen. 2025. WMNav: Integrating Vision-Language Models into World Models for Object Goal Navigation. https://api.semanticscholar.org/CorpusID:276776282
- Pal et al. (2021) Anwesan Pal, Yiding Qiu, and Henrik Christensen. 2021. Learning hierarchical relationships for object-goal navigation. In Conference on Robot Learning. PMLR, 517–528.
- Ramakrishnan et al. (2022) Santhosh Kumar Ramakrishnan, Devendra Singh Chaplot, Ziad Al-Halah, Jitendra Malik, and Kristen Grauman. 2022. Poni: Potential functions for objectgoal navigation with interaction-free learning. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 18890–18900.
- Shah et al. (2022) Dhruv Shah, Blazej Osinski, Brian Ichter, and Sergey Levine. 2022. LM-Nav: Robotic Navigation with Large Pre-Trained Models of Language, Vision, and Action. In 6th Annual Conference on Robot Learning. https://openreview.net/forum?id=UW5A3SweAH
- Song et al. (2025) Daeun Song, Jing Liang, Amirreza Payandeh, Amir Hossain Raj, Xuesu Xiao, and Dinesh Manocha. 2025. VLM-Social-Nav: Socially Aware Robot Navigation Through Scoring Using Vision-Language Models. IEEE Robotics and Automation Letters 10, 1 (2025), 508–515. doi:10.1109/LRA.2024.3511409
- Wang et al. (2023) Hongcheng Wang, Andy Guan Hong Chen, Xiaoqi Li, Mingdong Wu, and Hao Dong. 2023. Find what you want: Learning demand-conditioned object attribute space for demand-driven navigation. Advances in Neural Information Processing Systems 36 (2023), 16353–16366.
- Wang et al. (2024c) Hongcheng Wang, Peiqi Liu, Wenzhe Cai, Mingdong Wu, Zhengyu Qian, and Hao Dong. 2024c. MO-DDN: A Coarse-to-Fine Attribute-based Exploration Agent for Multi-object Demand-driven Navigation. In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
- Wang et al. (2024a) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, et al. 2024a. Qwen2-vl: Enhancing vision-language model’s perception of the world at any resolution. arXiv preprint arXiv:2409.12191 (2024).
- Wang et al. (2024b) Tian Wang, Junming Fan, and Pai Zheng. 2024b. An LLM-based vision and language cobot navigation approach for Human-centric Smart Manufacturing. Journal of Manufacturing Systems 75 (Aug. 2024), 299–305. doi:10.1016/j.jmsy.2024.04.020 Publisher Copyright: © 2024 The Society of Manufacturing Engineers.
- Wang et al. (2024d) Xudong Wang, Weihong Ren, Xi’ai Chen, Huijie Fan, Yandong Tang, and Zhi Han. 2024d. Uni-YOLO: Vision-Language Model-Guided YOLO for Robust and Fast Universal Detection in the Open World. In Proceedings of the 32nd ACM International Conference on Multimedia. 1991–2000.
- Wang et al. (2018) Xin Wang, Wenhan Xiong, Hongmin Wang, and William Yang Wang. 2018. Look before you leap: Bridging model-free and model-based reinforcement learning for planned-ahead vision-and-language navigation. In Proceedings of the European Conference on Computer Vision (ECCV). 37–53.
- Wason and Evans (1974) Peter C Wason and J St BT Evans. 1974. Dual processes in reasoning? Cognition 3, 2 (1974), 141–154.
- Wei et al. (2024) Jiateng Wei, Quan Lu, Ning Jiang, Siqi Li, Jingyang Xiang, Jun Chen, and Yong Liu. 2024. Structured Optimal Brain Pruning for Large Language Models. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing. 13991–14007.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35 (2022), 24824–24837.
- Wu et al. (2024) Pengying Wu, Yao Mu, Bingxian Wu, Yi Hou, Ji Ma, Shanghang Zhang, and Chang Liu. 2024. Voronav: Voronoi-based zero-shot object navigation with large language model. arXiv preprint arXiv:2401.02695 (2024).
- Xiang and Zhang (2024) Jingyang Xiang and Sai Qian Zhang. 2024. DFRot: Achieving Outlier-Free and Massive Activation-Free for Rotated LLMs with Refined Rotation. arXiv preprint arXiv:2412.00648 (2024).
- Yin et al. (2024) Zecheng Yin, † ChonghaoCheng, Yinghong Liao, Zhihao Yuan, Shuguang Cui, and Zhen Li. 2024. Navigation with VLM framework: Go to Any Language. ArXiv abs/2410.02787 (2024). https://api.semanticscholar.org/CorpusID:273163041
- Yue et al. (2024) Lu Yue, Dongliang Zhou, Liang Xie, Feitian Zhang, Ye Yan, and Erwei Yin. 2024. Safe-vln: Collision avoidance for vision-and-language navigation of autonomous robots operating in continuous environments. IEEE Robotics and Automation Letters (2024).
- Zhan et al. (2024) Zhaohuan Zhan, Lisha Yu, Sijie Yu, and Guang Tan. 2024. Mc-gpt: Empowering vision-and-language navigation with memory map and reasoning chains. arXiv preprint arXiv:2405.10620 (2024).
- Zhang et al. (2024b) Jiazhao Zhang, Kunyu Wang, Rongtao Xu, Gengze Zhou, Yicong Hong, Xiaomeng Fang, Qi Wu, Zhizheng Zhang, and He Wang. 2024b. Navid: Video-based vlm plans the next step for vision-and-language navigation. arXiv preprint arXiv:2402.15852 (2024).
- Zhang et al. (2025) Lingfeng Zhang, Xiaoshuai Hao, Qinwen Xu, Qiang Zhang, Xinyao Zhang, Pengwei Wang, Jing Zhang, Zhongyuan Wang, Shanghang Zhang, and Renjing Xu. 2025. MapNav: A Novel Memory Representation via Annotated Semantic Maps for VLM-based Vision-and-Language Navigation. ArXiv abs/2502.13451 (2025). https://api.semanticscholar.org/CorpusID:276449634
- Zhang et al. (2024a) Yue Zhang, Ziqiao Ma, Jialu Li, Yanyuan Qiao, Zun Wang, Joyce Chai, Qi Wu, Mohit Bansal, and Parisa Kordjamshidi. 2024a. Vision-and-Language Navigation Today and Tomorrow: A Survey in the Era of Foundation Models. ArXiv abs/2407.07035 (2024). https://api.semanticscholar.org/CorpusID:271064503
- Zhang et al. (2024c) Ying Zhang, Peng Zhang, Mincong Huang, Jingyang Xiang, Yujie Wang, Chao Wang, Yineng Zhang, Lei Yu, Chuan Liu, and Wei Lin. 2024c. Qqq: Quality quattuor-bit quantization for large language models. arXiv preprint arXiv:2406.09904 (2024).
- Zhao et al. (2023) Qianfan Zhao, Lu Zhang, Bin He, Hong Qiao, and Zhiyong Liu. 2023. Zero-shot object goal visual navigation. In 2023 IEEE International Conference on Robotics and Automation (ICRA). IEEE, 2025–2031.
- Zhou et al. (2024b) Gengze Zhou, Yicong Hong, Zun Wang, Xin Eric Wang, and Qi Wu. 2024b. Navgpt-2: Unleashing navigational reasoning capability for large vision-language models. In European Conference on Computer Vision. Springer, 260–278.
- Zhou et al. (2024a) Gengze Zhou, Yicong Hong, and Qi Wu. 2024a. Navgpt: Explicit reasoning in vision-and-language navigation with large language models. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 7641–7649.
A case of Demand Matching dataset
Appendix A Dataset Annotations
In this section, we present examples of datasets utilized to train the 3D Robot Perception, Demand Matching and Heuristic Process.
A.1. Annotations for 3D Robot Perception
We utilized data collected from the AI2THOR simulator to train the 3D Robot Perception Module. The process of obtaining 3D bounding box information involved generating point clouds for each object using segmentation and depth maps provided by the AI2THOR simulator. These point clouds were used to extract the minimal bounding box that fully encloses each object, allowing us to derive accurate 3D bounding box information.
Each frame in the dataset contains the following detailed annotations:
bbox2D_tight: The 2D corners of the annotated tight bounding box, tightly fitting the object in 2D image space.
bbox2D_proj: The 2D corners projected from the 3D bounding box, representing the 2D location of the object from the camera’s perspective.
bbox3D_cam: The 3D coordinates (in meters) of the eight corners of the bounding box in camera coordinates.
center_cam: The 3D coordinates of the object’s center in camera space (in meters).
dimensions: The 3D dimensions of the object in meters (width, height, and length).
R_cam: A 3x3 rotation matrix that represents the transformation between the object’s coordinate frame and the camera’s coordinate frame.
A.2. Annotations for Demand Matching
In this study, we curated a specialized instruction-attribute dataset to train the Demand Matching Module. The dataset construction process involved leveraging the provided instruction data, which was paired with the objects that could potentially fulfill the instructions.
The process begins with each instruction and its associated objects being input into a Large Language Model (LLM), which generates the common attributes shared by objects that could satisfy the instruction. These attributes are crucial for matching objects with the instruction. A QA-format dataset is then created, where the first part consists of the common attributes that define the correct object, assisting the Demand Matching Module in identifying the appropriate object. The second part includes a list of objects, selected from a predefined set, that align with the generated attributes and are deemed to fulfill the instruction. A case of Demand Matching dataset is shown in Figure 8.
A.3. Annotations for Heuristic Process
To support the supervised fine-tuning of our Heuristic Process, we created datasets using the AI2THOR simulator. For each scenario, we randomly select an instruction and target object, then compute an optimal path using the A* algorithm, which is discretized into a sequence of actions guiding the agent to the target. After each action, the simulator produces updated images, used to identify the target object, enabling precise localization. When the agent detects the target, VLM generates a structured scene description and reasoning based on the target’s spatial location and navigation decisions.
As shown in Figure 9, the data is formatted as a Visual Question Answering (VQA) structure, with a question about the information of target object and an answer containing the scene description, reasoning steps, and final decision. The structured data is stored in the knowledge base, allowing the system to build a robust understanding of the scene and make informed decisions in future scenarios.
A case of knowledge base dataset
Appendix B Prompt Details
We provide a detailed description of the demand matching prompt in Figure 10, which is used to align the object in the field of view with the given instruction. Besides, Figure 11 illustrates the explore prompt employed by the Heuristic Process for exploration when no target object is present in the field of view. Additionally, Figure 13 depicts the exploit prompt utilized by the Heuristic Process, which guides the approach toward the matched objects with precision and accuracy. Furthermore, Figure 13 presents the reflection prompt, employed by the Analytic Process, which incorporates key elements from the previous prompts and includes criteria for detecting potential errors in the preceding frame. Moreover, Figure 15 highlights the attitude generation and training prompt. Finally, as depicted in Figure 15, the A* algorithm is employed to generate the Description and Reasoning prompts, which explain the selected path and the logical steps behind the decision-making process.
System prompt for Demand Matching
System prompt for Explore Module
System prompt for Exploit Module
System prompt for reflection mechanism
System prompt about the generation and train of attribute
System prompt about the generation of scene description and reasoning
A case of Explore module
A straight case in an Exploit module
A bypass case in an Exploit module
Case study for reflection mechanism.
Appendix C Visualization Cases
C.1. Cases of Explore
Figure 17 presents a case to show the results using the Explore. When no object in the field of view matches the instruction, the VLMs take images of the current scene, historical actions, and recent rotations, generating scene descriptions, reasoning, and decisions, which include a series of actions to explore. As shown in this case, the historical actions included ’RotateLeft’, and the recent rotation data revealed that the last two rotations were also ’RotateLeft’ (consecutive rotations in one direction are recorded once). After passing through a door, a large open area was detected ahead. The VLM correctly decided on a series of ’MoveAhead’ actions to explore forward while avoiding redundant exploration.
C.2. Cases of Exploit
We also provide several examples to demonstrate the qualitative performance of our proposed CogDDN, as illustrated in Figure 17 and 18. When an object matching the instruction is present in the field of view, the VLMs capture images of the current scene, generate scene descriptions and reasoning, and produce a decision that includes a single action to approach the target object. As shown in 17, the VLM correctly calculated the angle between the toilet and the agent, detected that there were no obstacles on the path to the toilet, and ultimately generated a ’MoveAhead’ decision to approach it. Figure 18 presents a challenging scenario where the target object, a painting, is directly ahead, but a bed on the floor obstructs the agent’s forward movement. The system effectively perceives the obstacle ahead, analyzes the environment, and promptly navigates around it towards a feasible area. In this case, with a wall on the right and a large feasible area on the left, the system correctly made a ’RotateLeft’ decision, demonstrating its ability to effectively plan and adapt to the surroundings.
C.3. Cases of reflection mechanism
We also present an example in Figure 19 to illustrate the reflection mechanism. When the current path encounters obstacles, the image, description, reasoning, and decision are input into the Analytic Process to detect potential errors and provide reflective reasoning and decisions. As shown in this case, the initial analysis by the Heuristic Process incorrectly identifies the feasible area ahead, which requires immediate attention. In contrast, during the reflection phase, the Analytic Process accurately identifies the chair as the primary obstacle. It reassesses the previous ’MoveAhead’ decision, suggesting that it may have been based on an inference error, as rotation would have been necessary to avoid colliding with the chair in this situation.
Appendix D Implementation Details
We train the UniMODE model with a batch size of 64 for 30,000 iterations on 4 RTX6000 GPUs, taking approximately 10 hours. The input image resolution is 448 × 448 pixels, resized to 640 × 512 pixels. We use the AdamW optimizer with a learning rate of , weight decay of , and a CosineAnnealing scheduler with a warm-up factor of 0.3333 for the first 2000 iterations. Gradient clipping is applied at 35.0 for stability.
For supervised fine-tuning (SFT), we use the Qwen2-VL-7B model with LoRA, training with a batch size of 1 for five epochs on 4 RTX6000 GPUs in 1 hour. Key hyperparameters include a learning rate of , gradient accumulation of 8, and a cosine scheduler. LoRA is applied with a rank of 64 and alpha of 128, while freezing the ViT layers.
For the Heuristic Process, we conduct SFT on Qwen2-VL-7B for five epochs using knowledge base samples, which takes about 14 hours, with consistent hyperparameters from the Demand Matching module.

### v2

@@
 1. Introduction
-Mobile robots are increasingly essential for assisting humans in unknown and unstructured environments, such as households, hospitals, and warehouses.
-To operate effectively, they must be capable of interpreting and responding to implicit human demands, even when target objects are uncertain or not explicitly specified.
-For instance, a hungry person instinctively seeks food based on internal cues rather than direct instructions.
-Similarly, demand-driven navigation (DDN) (Wang et al., 2023, 2024c) enables robots to locate objects that fulfill such implicit needs without relying on predefined object lists or locations.
-However, as shown in Figure 1 (a), conventional data-driven DDN methods (Wang et al., 2023, 2024c) depend heavily on extensive pre-collected data for model training and decision-making, limiting their generalization to novel scenarios and vague instructions.
+Mobile robots are becoming indispensable in dynamic and unstructured environments such as households, hospitals, and warehouses.
+To operate effectively and naturally within these human-centric spaces, they must move beyond simple command execution to interpret and respond to implicit human demands.
+This capability is particularly critical when target objects are uncertain or not explicitly specified.
+For instance, a hungry person instinctively seeks food based on internal cues, rather than following a direct instruction like ”find the apple on the kitchen counter”.
+Similarly, demand-driven navigation (DDN) (Wang et al., 2023, 2024c) enables robots to locate objects that fulfill such implicit needs without relying on predefined object lists or known locations.
+However, as illustrated in Figure 1 (a), conventional data-driven DDN methods (Wang et al., 2023, 2024c) are fundamentally constrained by their heavy reliance on extensive, pre-collected datasets for training and decision-making.
+This dependence curtails their adaptability and generalization when encountering novel scenarios or vaguely phrased instructions.
 Comparison between cognition-driven and data-driven.
-Recent breakthroughs in Large Language Models (LLMs) and Vision-Language Models (VLMs) have demonstrated strong reasoning capabilities and extensive world knowledge, making them well-suited for navigation tasks.
-Cognition-driven approaches (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025; Wang et al., 2024b; Shah et al., 2022) leverage these models to help agents interpret instructions and locate target objects.
-The DDN tasks combine natural language with visual perception, requiring agents to infer and act upon implicit user intent.
-As illustrated in Figure 1 (b), our framework integrates LLMs and VLMs for DDN, enabling more sophisticated reasoning over vague or ambiguous instructions.
-Additionally, VLMs improve navigation by jointly processing visual and linguistic inputs, effectively bridging semantic intent with spatial context.
-This cognition-driven, multi-agent approach enhances the agent’s spatial and contextual understanding, resulting in more accurate and adaptable navigation.
-Figure 1 (c) shows that the cognition-driven method generalizes better than traditional data-driven approaches, exhibiting a smaller performance drop between seen and unseen scenarios.
+Recent breakthroughs in Large Language Models (LLMs) and Vision-Language Models (VLMs) have catalyzed a new era in robotic intelligence, as these models possess remarkable reasoning capabilities and extensive world knowledge that are exceptionally well-suited for complex navigation tasks.
+Leveraging these advancements, cognition-driven approaches (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025; Wang et al., 2024b; Shah et al., 2022) employ these models to help agents interpret high-level instructions and strategically locate targets.
+The DDN task, which inherently fuses natural language with real-time visual perception, demands that an agent infer and act upon subtle user intent.
+As depicted in Figure 1 (b), our proposed framework synergistically integrates LLMs and VLMs for DDN, enabling sophisticated and nuanced reasoning over ambiguous instructions.
+The embedded VLMs enhance navigation by jointly processing visual and linguistic inputs, effectively bridging the semantic gap between high-level intent and spatial context.
+This cognition-driven, multi-agent approach significantly improves the agent’s spatial and contextual understanding, yielding more accurate, robust, and adaptable navigation.
+Figure 1 (c) shows that the cognition-driven method generalizes better than traditional data-driven approaches, exhibiting a substantially smaller performance drop when transitioning from seen to unseen scenarios.
 CogDDN Framework for Demand-Driven Navigation Tasks.
-This paper presents CogDDN, a novel VLM-based approach inspired by the dual-process framework for closed-loop decision-making.
-A key distinction is that it emulates human cognitive and learning mechanisms.
-Unlike many VLM-based navigation systems (Long et al., 2024; Cai et al., 2024) that rely on additional views, depth, or maps, our system uses only front-facing views, aligning with human search behavior and real-world robotic constraints.
-At a result, our system performs a more realistic navigation task by mirroring the perceptual constraints and decision-making processes of human object search.
-In parallel with human decision-making, CogDDN incorporates a demand-matching module that leverages LLMs to accurately link instructions to relevant objects, allowing the agent to identify and locate key items that satisfy implicit user needs.
-Furthermore, we introduce a dual-process decision-making system based on dual-process theory (Kahneman, 2011; Evans and Stanovich, 2013; Wason and Evans, 1974), which simulates human cognition through two components: a Heuristic Process (System-I) and an Analytic Process (System-II) (Mei et al., 2024; Ma et al., 2025).
-The Heuristic Process enables rapid, intuitive decisions based on prior knowledge, while the Analytic Process refines these decisions by analyzing new data and updating the knowledge base.
-Before navigation, we construct a knowledge base to store high-quality strategies adaptable to diverse scenarios.
-The Heuristic Process utilizes this accumulated knowledge, refined through supervised fine-tuning (SFT), to guide initial decisions.
-Meanwhile, the Analytic Process updates the knowledge base when encountering obstacles or challenges, enabling continuous learning and improved decision-making.
-In contrast to direct end-to-end methods, our system integrates Chain of Thought (CoT) (Wei et al., 2022) reasoning, enhancing the agent’s ability to reason through complex scenarios and make informed decisions, rather than merely outputting final actions.
-The main contributions are summarized as follows:
-We propose CogDDN, a novel VLM-based framework for DDN tasks that simulates human-like cognitive mechanisms and learning processes.
-We introduce a dual-process decision-making module.
-Without human involvement, our approach allows the empirical Heuristic Process to inherit the capabilities of the rational Analytic Process through a self-supervised mechanism.
-CogDDN uses the Analytic Process and a knowledge accumulation mechanism to enrich a transferable knowledge base, enabling continual learning and effective generalization to novel navigation environments.
+This paper introduces CogDDN, a novel VLM-based framework for DDN that is inspired by the human dual-process cognitive framework and designed for robust, closed-loop decision-making.
+A key differentiator of our approach is its emulation of human cognitive and learning mechanisms to solve complex search tasks.
+Unlike many contemporary VLM-based navigation systems (Long et al., 2024; Cai et al., 2024) that rely on supplementary data such as multiple camera views, depth information, or pre-built maps, our system uses only egocentric, front-facing visual input.
+This design choice intentionally aligns our agent’s perception with natural human search behavior and adheres to the practical constraints of real-world robotic hardware.
+Consequently, our system addresses a more realistic and challenging navigation task by mirroring the perceptual limitations and sequential decision-making inherent in human object search.
+Analogous to human cognitive processes, CogDDN incorporates a demand-matching module that leverages the semantic reasoning power of LLMs.
+This module translates ambiguous natural language instructions into a set of relevant, potential target objects, enabling the agent to identify and prioritize items that satisfy the user’s underlying needs.
+Furthermore, we introduce a core decision-making system founded on dual-process theory (Kahneman, 2011; Evans and Stanovich, 2013; Wason and Evans, 1974), which simulates human cognition via two distinct yet complementary components: a Heuristic Process (System-I) and an Analytic Process (System-II) (Mei et al., 2024; Ma et al., 2025).
+The Heuristic Process, akin to human intuition, facilitates rapid and efficient decisions based on prior knowledge.
+In contrast, the Analytic Process employs deliberate, step-by-step reasoning to refine these decisions, particularly in novel or challenging situations, and updates the knowledge base accordingly.
+Prior to navigation, we construct a knowledge base of high-quality, generalizable strategies.
+The Heuristic Process, continuously refined through supervised fine-tuning (SFT), draws upon this accumulated knowledge to guide initial actions.
+Meanwhile, the Analytic Process is invoked to update this knowledge base upon encountering obstacles or failures, creating a virtuous cycle of continuous learning and performance improvement.
+In contrast to direct end-to-end methods, our system integrates Chain of Thought (CoT) (Wei et al., 2022) reasoning, enhancing the agent’s ability to logically reason through complex scenarios and make transparent, informed decisions, rather than merely outputting final actions.
+The main contributions of this paper are summarized as follows:
+We propose CogDDN, a novel VLM-based framework for DDN tasks that simulates human-like cognitive mechanisms and learning processes to navigate based on implicit demands.
+We introduce a dual-process decision-making module wherein the empirical Heuristic Process learns from the rational Analytic Process via a self-supervised mechanism, eliminating the need for human intervention.
+CogDDN uses the Analytic Process and a knowledge accumulation mechanism to progressively enrich a transferable knowledge base, enabling continual learning and effective generalization to novel navigation environments.
 2. Related Works
 2.1. Demand-driven Navigation
-Demand-driven Navigation (DDN) (Wang et al., 2023, 2024c) is an innovative task requiring an agent to navigate an environment based on high-level human demands described in natural language.
-Unlike traditional visual navigation tasks like Visual Object Navigation (VON) (Ramakrishnan et al., 2022; Zhao et al., 2023; Kiran et al., 2022; Chaplot et al., 2020; Gadre et al., 2022; Pal et al., 2021) and Visual Language Navigation (VLN) (Wu et al., 2024; Yue et al., 2024; Zhang et al., 2024b; Zhou et al., 2024a, b; Hong et al., 2025; Goetting et al., 2024; Cai et al., 2024), DDN does not focus on specific object categories or step-by-step instructions but rather on fulfilling user-defined needs.
-Using abstract, high-level instructions, DDN enables the agent to infer which objects in the environment are relevant to the given demand.
-This makes DDN more flexible and dynamic than VON or VLN, as it centers on addressing ”human demands” rather than specific objects.
-The agent must interpret these demands and make inferences from the current scene, providing a more generalized approach to navigation.
-The main challenge of DDN lies in interpreting and fulfilling human demand within the visual context.
+Demand-driven Navigation (DDN) (Wang et al., 2023, 2024c) represents an advanced navigation paradigm where an agent must interpret and act upon high-level human demands expressed in natural language.
+This task diverges significantly from traditional visual navigation frameworks such as Visual Object Navigation (VON) (Ramakrishnan et al., 2022; Zhao et al., 2023; Kiran et al., 2022; Chaplot et al., 2020; Gadre et al., 2022; Pal et al., 2021), which typically requires finding a specific object category (e.g., ”find a chair”), and Visual Language Navigation (VLN) (Wu et al., 2024; Yue et al., 2024; Zhang et al., 2024b; Zhou et al., 2024a, b; Hong et al., 2025; Goetting et al., 2024; Cai et al., 2024), which relies on explicit, step-by-step instructions (e.g., ”walk past the table and turn left”).
+In contrast, DDN centers on fulfilling an underlying user need (e.g., ”I’m thirsty”), compelling the agent to infer which objects in the environment are relevant to satisfying the stated demand.
+This formulation makes DDN a more flexible and dynamic, shifting the focus from locating specific objects to addressing abstract human demands.
+The agent is thus required to perform a higher level of semantic interpretation and contextual inference based on the current visual scene, fostering a more generalized and human-aligned approach to navigation.
+Therefore, the primary challenge in DDN lies in this sophisticated process of interpreting abstract human needs and grounding them within a rich, dynamic visual context to guide action.
 2.2. Transitioning from Data-Driven to Cognition-Driven Navigation
-Early navigation methods were predominantly data-driven, relying on imitation learning (Nguyen et al., 2019; Kamath et al., 2023) and reinforcement learning (Wang et al., 2018; Chen et al., 2022) to train reactive policies on large-scale labeled datasets.
-While these approaches achieved promising results within their training distributions, they often failed to generalize to out-of-distribution scenarios (Cai et al., 2025; Zhan et al., 2024), mainly due to limited semantic reasoning and contextual awareness, and often incur high computational costs during both training and deployment (Zhang et al., 2024c; Wei et al., 2024; Xiang and Zhang, 2024).
-Motivated by the success of pre-trained vision-language models in various multimodal reasoning tasks (Ma et al., 2024; Wang et al., 2024d; Lu et al., 2025; Liu et al., 2025a, b), recent navigation research has begun to adopt these models as cognitive priors to move toward more robust and human-like navigation strategies (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025).
-Inspired by human-like reasoning, these systems unify perception and language to support language-driven goal specification (Mirjalili et al., 2024) and semantic understanding (Bangguo Yu and Cao, 2024).
-Unlike traditional methods that depend on structured instructions or spatial maps, VLM-based approaches allow agents to interpret open-ended natural language commands and navigate in complex, unfamiliar environments (Zhang et al., 2024a; Song et al., 2025).
-These systems emphasize multimodal grounding, context-aware planning, and continual learning, and they often incorporate reasoning techniques such as Chain-of-Thought to support interpretable and generalizable behavior.
+Early approaches to robotic navigation were predominantly data-driven, relying heavily on imitation learning (Nguyen et al., 2019; Kamath et al., 2023) and reinforcement learning (Wang et al., 2018; Chen et al., 2022) to train reactive policies.
+These methods learn to map sensory inputs directly to actions by training on large-scale, often manually labeled, datasets.
+While such approaches have demonstrated promising results within their training distributions, they frequently exhibit brittleness and fail to generalize to out-of-distribution scenarios (Cai et al., 2025; Zhan et al., 2024).
+This limitation arises primarily from a lack of deep semantic reasoning and contextual awareness, as the models often learn spurious correlations in the data rather than robust, causal relationships.
+Furthermore, these data-intensive methods typically incur high computational costs during extensive training and deployment (Zhang et al., 2024c; Wei et al., 2024; Xiang and Zhang, 2024).
+Motivated by the success of pre-trained vision-language models in a wide array of multi-modal reasoning tasks (Ma et al., 2024; Wang et al., 2024d; Lu et al., 2025; Liu et al., 2025a, b), recent navigation research has shifted towards cognition-driven paradigms.
+This new wave of research adopts large-scale models as powerful cognitive priors, aiming to instill more robust and human-like navigation strategies (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025).
+Inspired by human reasoning, these systems unify perception, action, and language to support flexible, language-driven goal specification (Mirjalili et al., 2024) and a deeper semantic understanding of the world (Bangguo Yu and Cao, 2024).
+Unlike traditional methods that depend on structured instructions or explicit spatial maps, VLM-based approaches empower agents to interpret open-ended natural language commands and navigate effectively in complex, unfamiliar environments (Zhang et al., 2024a; Song et al., 2025).
+These modern systems emphasize multi-modal grounding, context-aware planning, and continual learning.
+To enhance transparency and robustness, they often incorporate explicit reasoning techniques, such as Chain-of-Thought (CoT) prompting, which encourages the model to generate a sequence of logical steps, thereby supporting more interpretable and generalizable navigation behavior.
 A detailed architecture diagram of the proposed CogDDN.
@@
 3.1. Task Definition
-CogDDN tasks require agents to locate relevant objects in a given environment based on human demands and perceptual inputs.
-They can be formalized as a Partially Observable Markov Decision Process (POMDP), defined as
+CogDDN tasks require agents to locate the relevant objects in a given environment based on human demands and perceptual inputs.
+These tasks can be formally described as a Partially Observable Markov Decision Process (POMDP), defined as
 where represents the state space (the agent’s current state), is the observation space (including human demands and images), is the action space (e.g., MoveAhead, RotateLeft, LookDown, etc.), and is the action execution function.
-For instance, a user may state, ”I am thirsty.” In response, the agent must identify an object capable of quenching thirst based on the current visual observation and then execute the actions to locate the object.
+For instance, a user may state, ”I am thirsty”.
+In response, the agent must identify an object capable of quenching thirst based on the current visual observation and then execute the actions to locate the object.
 After each action, the agent receives updated observations and generates new actions based on the updated information to continue searching for the object.
-The task is complete once the agent is within 1.5 meters of the target object.
+The task is considered complete when the distance between the agent and the target object is reduced to within 1.5 meters.
 3.2. Demand Matching
-Objects that fulfill the same human demand in demand-driven navigation share key properties.
-For instance, paintings, houseplants, and statues are suitable candidates for decorating a space, as they enhance visual appeal and complement the intended environment.
-This relationship between human demands and object properties is grounded in universal knowledge.
-While LLMs are proficient in reasoning based on instructions and properties, they encounter difficulties when an exact match for a request is unavailable.
-In such cases, LLMs may suggest suboptimal objects.
-For instance, if a user requests, ”I need something to hold my flowers,” a mug might be suggested despite being a less-than-ideal choice.
-This tendency significantly reduces the model’s accuracy in selecting the most suitable object.
-To mitigate this issue, we apply supervised fine-tuning (SFT) to train the LLM, allowing it better to align objects with user requirements.
-This method improves the model’s capacity to identify and recommend the most appropriate objects, ensuring it avoids suggesting less optimal options, even when an exact match is not readily apparent.
-Formally, the Demand Matching module implements a target objects generation function , where represents the space of observation space including human demands and images, is the space of demand properties, and is the space of target objects.
-By leveraging a fine-tuned LLM for demand matching, the system benefits from its ability to reason over complex instructions and object properties.
-This capability prevents it from forcing matches between human demands and objects when an exact match is unavailable, ensuring more appropriate suggestions even in ambiguous scenarios.
+A foundational principle of demand-driven navigation is that objects capable of fulfilling the same high-level human demand often share a set of key, underlying properties.
+For instance, paintings, houseplants, and statues are all suitable candidates for decorating a space.
+Although physically distinct, they all enhance visual appeal and complement the ambient environment.
+This intrinsic relationship between abstract human demands and object properties is grounded in commonsense or universal knowledge.
+While LLMs are proficient at general reasoning based on instructions and properties, they can encounter difficulties when a perfect match for a given request is unavailable in the immediate environment.
+In such cases, standard LLMs may suggest suboptimal or functionally adjacent objects.
+For example, if a user requests, ”I need something to hold my flowers”, and no vase is present, the model might suggest a mug.
+Although a mug can technically hold water and flowers, it is a less-than-ideal choice domestically.
+This tendency to over-generalize based on basic affordances can significantly reduce the model’s accuracy in selecting the most suitable object that aligns with user expectations.
+To mitigate this issue, we employ supervised fine-tuning (SFT) to further train the LLM, allowing it to better align object affordances with nuanced user requirements and contextual appropriateness.
+The fine-tuning enhances the model’s capacity to identify and recommend the most fitting objects, ensuring it avoids suggesting less optimal options, even when an exact match is not readily apparent.
+Formally, the Demand Matching module implements a target objects generation function , where represents the space of observation space including the human demands and visual information from images, is the latent space of inferred demand properties, and is the resulting space of candidate target objects.
+By leveraging a fine-tuned LLM for demand matching, the system benefits from an enhanced ability to reason over complex instructions and subtle object properties.
+This specialized capability prevents it from forcing matches between human demands and available objects when a direct correspondence is absent, ensuring more appropriate suggestions even in ambiguous scenarios.
 Please refer to Appendix A.2 for annotation details.
@@
 2) Reasoning: This component leverages the visual-language model’s spatial awareness and commonsense reasoning capabilities based on the target’s location and the scene description to guide the agent’s navigation.
-3) Decision: Drawing from the output of the reasoning component, the model generates the optimal decision for the current navigation task, ensuring the selection of the most effective action.
+3) Decision: Drawing from the output of the reasoning component, the model generates the optimal decision for the current navigation task, ensuring the most effective action.
 This process incorporates knowledge transfer mechanisms to enhance navigation performance.
 Building on this foundational approach of leveraging scene descriptions and reasoning for navigation optimization, we construct a knowledge base that integrates environmental data with dynamic decision-making processes.
-As illustrated in Figure 4, during the execution of the trajectory generated by the A* algorithm (Hart et al., 1968), once the target object is detected, the current viewpoint image , target object , and final action S are input into the VLM to generate scene descriptions D and reasoning R, which are then recorded as an experience in the knowledge base.
+As illustrated in Figure 4, during the execution of the trajectory generated by the A* algorithm (Hart et al., 1968), once the target object is detected, the current viewpoint image , target object , and final action S are input into the VLMs to generate scene descriptions D and reasoning R, which are then recorded as an experience in the knowledge base.
 This accumulated experience can be progressively transferred to the Heuristic Process.
@@
 3.4. Heuristic Process
-While the VLM effectively generates reasoning based on decisions from the A* algorithm (Hart et al., 1968), it struggles to directly infer the relationship between reasoning and decisions.
-Moreover, its relatively slow processing speed often leads to redundancy and repetition, limiting practical use in indoor navigation.
-Formally, the Heuristic Process module implements an action execution function , where represents the space of target objects, depicts the state space, and is the action space, which is presented in table 1.
-Inspired by human navigation, where efficiency is gained through repeated practice and environmental familiarity, allowing individuals to navigate with minimal cognitive effort, we introduce the Heuristic Process in CogDDN.
+While the VLMs can effectively generate reasoning that aligns with decisions from path planning algorithms like A* algorithm (Hart et al., 1968), they often struggle to directly infer the causal relationship between reasoning and decisions.
+Moreover, their relatively slow processing speed can lead to redundancy and repetition, limiting their practical application in real-time indoor navigation.
+To address these challenges, we introduce the Heuristic Process in CogDDN.
+This process is inspired by human navigation, where efficiency is gained through repeated practice and environmental familiarity, eventually allowing individuals to navigate with minimal cognitive effort.
+Formally, the Heuristic Process module implements an action execution function , where represents the space of target objects, depicts the state space, and is the action space, as presented in Table 1.
 As demonstrated in Figure 5, it consists of two modules: Explore and Exploit, both leveraging Chain-of-Thought (CoT) reasoning to adapt to new environments and optimize learned navigation execution.
@@
 Explore.
-The Explore module is activated when the Demand Matching module cannot identify a suitable target within the current scene.
-In such cases, the system searches the environment by generating exploratory actions based on its visual perspective, utilizing the VLM.
-This aims to uncover previously overlooked objects or paths.
-As Figure 2 shows, the system first generates a Scene Description D based on the current viewpoint image .
-Subsequently, it conducts Reasoning R to direct the exploration process.
-The system’s Decision S consists of a series of actions to explore unvisited regions.
-The module incorporates information from prior actions and recent rotations to minimize redundant exploration.
-If obstacles hinder progress, the flag is set to true, prompting the system to dynamically generate new exploratory actions, ensuring continuous progress without unnecessary repetition or resource depletion.
+The Explore module is activated when the Demand Matching module fails to identify a suitable target object within the agent’s current field of view.
+In this mode, the system defaults to a broad, information-seeking search of the environment.
+It generates exploratory actions by leveraging the VLM to interpret its visual perspective and reason about promising, unvisited areas.
+The objective is to efficiently survey the surroundings to uncover previously overlooked objects or navigational paths.
+As illustrated in Figure 2, this process begins with the system generating a concise Scene Description D from the current viewpoint image .
+Subsequently, it engages in explicit Reasoning R to formulate a strategy for exploration.
+Based on this reasoning, the system makes a Decision S, which comprises a sequence of actions designed to navigate toward unvisited regions.
+To prevent inefficient behavior, the module incorporates a short-term memory of prior actions and recent rotations to minimize redundant movements.
+Furthermore, if an obstacle impedes progress, a flag is set to true, prompting the system to dynamically replan and generate new exploratory actions, which ensures continuous progress without getting stuck in loops or depleting resources.
 | Action | Definition |
@@
 Exploit.
-Once the Demand Matching module identifies a target object, the system transitions to the Exploit module.
-Unlike the Explore phase, which focuses on broader navigation and search strategies, the Exploit phase emphasizes precise, deliberate actions to achieve the navigation goal.
-This phase leverages a fine-tuned version of the VLM, optimized through prior exploration and experiences stored in the knowledge base.
-Fine-tuning improves the model’s ability to make accurate predictions and decisions based on subtle environmental cues, enabling the system to handle complex tasks with high precision.
-Guided by CoT, the system first generates a Scene Description D based on the current viewpoint image and the information of target objects , performs Reasoning R to infer the most suitable course of action based on the D and , and produces the most optimal Decision S, a single action focused on reaching the target.
-Focusing on one action rather than multiple potential actions improves efficiency and computational performance, ensuring rapid and effective goal achievement.
+Once the Demand Matching module successfully identifies a target object, the system transitions from exploration to exploitation.
+The Exploit module shifts the agent’s strategy from a broad, open-ended search to precise, goal-directed action.
+This phase leverages a version of the VLM that has been fine-tuned on prior successful experiences stored in the knowledge base, optimizing it for goal achievement.
+This fine-tuning enhances the model’s ability to make accurate predictions and decisions based on subtle environmental cues, enabling the system to execute complex navigation tasks with high precision.
+Guided by a Chain-of-Thought(CoT) process, the system first generates a Scene Description D based on the current viewpoint image and the information about the target objects .
+It then performs focused Reasoning R to infer the most direct course of action toward the target, given D and .
+Finally, it produces the most optimal Decision S, a single, high-confidence action focused on progressing toward the target.
+By concentrating on a single, decisive action rather than a sequence of potential actions, the system improves its decisiveness and computational efficiency, ensuring rapid and effective goal attainment.
+3.5. Analytic Process
+The Analytic Process serves as the deliberative reasoning component of our framework, designed to reflect upon and learn from navigation failures, as illustrated in Figure 6.
+Through comprehensive pre-training on diverse, large-scale datasets, VLMs naturally accumulate a vast repository of world knowledge.
+This enables them to address complex challenges with sophisticated reasoning and insight, which is a capability that aligns perfectly with the objectives of the Analytic Process.
+This process moves beyond the rapid, intuitive responses of the Heuristic Process and instead relies on detailed causal analysis and deep contextual understanding to draw sound, generalizable inferences from navigational errors encountered in indoor environments.
+Specifically, within the closed-loop operation of our system, any obstruction or failure encountered during the execution of the Heuristic Process triggers this reflective mechanism.
+When triggered, the Analytic Process utilizes the extensive world knowledge embedded within the VLMs to diagnose the situation.
+To do this, it ingests a snapshot of the state preceding the failure, including the target object’s location, the scene description, and the specific reasoning and decision that led to the error.
+The system then carefully analyzes the root cause of the event, identifies the flaw in the previous logic, and generates a corrected chain of reasoning along with a revised decision.
+The valuable insights gained from this reflective process, a successful pairing of a problem state with a corrected solution, are subsequently integrated into the knowledge base.
+This integration establishes an iterative learning loop, enabling the system to learn from its failures continuously and progressively enhance its decision-making capabilities.
+This virtuous cycle ultimately leads to more informed and accurate navigation strategies for future tasks and directly improves the performance of the Heuristic Process over time.
 Detailed procedure of the reflection mechanism.
 | Method | Seen Scene | Unseen Scene | ||||||||||
-| Seen Ins. | Unseen Ins. | Seen Ins. | Unseen Ins. | |||||||||
+| Seen Ins | Unseen Ins | Seen Ins | Unseen Ins | |||||||||
 | NSR | SPL | SSR | NSR | SPL | SSR | NSR | SPL | SSR | NSR | SPL | SSR | |
@@
 | FBE-MiniGPT-4 (Wang et al., 2023) | 3.5 | 3.0 | 2.2 | 3.5 | 3.2 | 2.0 | 3.5 | 3.5 | 2.0 | 4.0 | 4.0 | 3.5 |
-| FBE-GPT* (Wang et al., 2023) | 5 | 4.3 | 4.3 | 5.5 | 5.0 | 5.5 | 4.5 | 4.3 | 4.5 | 5.5 | 5.0 | 5.5 |
+| FBE-GPT* (Wang et al., 2023) | 5.0 | 4.3 | 4.3 | 5.5 | 5.0 | 5.5 | 4.5 | 4.3 | 4.5 | 5.5 | 5.0 | 5.5 |
 | GPT-3-Prompt* (Wang et al., 2023) | 0.3 | 0.01 | 0 | 0.3 | 0.01 | 0 | 0.3 | 0.01 | 0 | 0.3 | 0.01 | 0 |
@@
 | DDN (Wang et al., 2023) | 21.5 | 9.8 | 7.5 | 19.3 | 9.4 | 4.5 | 14.2 | 6.4 | 5.7 | 16.1 | 8.4 | 6.0 |
-| CogDDN(ours) | 38.3 | 17.2 | 29.8 | 37.5 | 18.0 | 28.6 | 33.3 | 16.4 | 27.1 | 34.5 | 17.1 | 27.5 |
-3.5. Analytic Process
-The Analytic Process is employed to reflect on navigation obstacles, as illustrated in Figure 6.
-Through comprehensive pre-training on diverse datasets, VLMs naturally accumulate broad world knowledge, enabling them to address complex challenges with sophisticated reasoning and insight.
-This capability aligns with the objectives of the Analytic Process, which relies on detailed analysis and contextual understanding to draw sound inferences in indoor navigation.
-Specifically, in a closed-loop indoor navigation scenario where the Demand Matching and Heuristic Process modules operate, any obstruction encountered during navigation will trigger the reflection mechanism.
-The Analytic Process utilizes the vast world knowledge embedded within the VLMs to analyze the situation based on previously gathered information.
-During this process, the object location, scene description, reasoning, and decision from the preceding frame before the obstruction are forwarded to the Analytic Process.
-The system then carefully analyzes the cause of the event, identifies the error, and generates corrected reasoning and decisions.
-The insights gained from this reflection process are subsequently integrated into the knowledge base, enabling the system to learn from failures continuously.
-This iterative learning process progressively enhances the system’s decision-making capabilities, resulting in more informed and accurate navigation strategies for future tasks.
+[…]
 4. Experiments
 4.1. Data preparation
-The experiments use the AI2Thor (Kolve et al., 2017) and the ProcThor dataset (Deitke et al., 2022).
-Following DDN (Wang et al., 2023), we evaluate 600 scenes (200 each from ProcThor’s train, validation, and test sets).
-These environments contain 109 distinct object categories, each capable of fulfilling the demand instructions.
-For instructions, we use the 1692 training, 241 validation, and 485 test instructions provided by DDN (Wang et al., 2023).
+Our experiments are conducted in the AI2Thor (Kolve et al., 2017) simulator using the ProcThor dataset (Deitke et al., 2022).
+Following the setup of DDN (Wang et al., 2023), we evaluate our model on 600 scenes (200 from each of ProcThor’s train, validation, and test splits), containing 109 object categories.
+We use the instruction sets provided by DDN (Wang et al., 2023), comprising 1692 training, 241 validation, and 485 test instructions.
 Data for 3D Robot Perception.
-We used data from the AI2Thor simulator to train the 3D Robot Perception Module.
-Specifically, we collected 55K frames, ensuring that 70% of the data was allocated for training, 10% for validation, and 20% for testing.
+For the 3D Robot Perception Module, we collected 55K frames from AI2THOR, split into 70% training, 10% validation, and 20% testing sets.
+Each frame includes annotations for object category, 2D bounding box, and 3D box attributes.
+The 3D box information was obtained by extracting the minimal bounding box from the object’s point cloud, which was generated using the simulator’s depth and segmentation maps.
 For clarity, we provide details and methods about data in Appendix A.1.
 Data for Demand Matching.
-By leveraging the instruction data, we curated an instruction-attribute dataset for the Demand Matching Module.
-The dataset was constructed by inputting instructions and associated objects into the LLM, which generated common attributes shared by objects fulfilling the instructions.
-We collected 7.2K data for training, 1.4K for validation, and 2.1K for testing.
-Additional details are provided in Appendix A.2.
+For the Demand Matching Module, we created a dataset of 10.7K instruction-attribute pairs.
+The construction process involved using LLMs to generate common attributes for objects associated with a given instruction.
+This was structured into a QA format, where the question (Q) was an instruction and a set of candidate objects, and the answer (A) was the shared attributes and a list of correct objects.
+The dataset was split into 7.2K training, 1.4K validation, and 2.1K test samples.
+Additional details are shown in Appendix A.2.
 Data for Heuristic Process.
-To support the supervised fine-tuning of our Heuristic Process using the knowledge base, we curated datasets by collecting data from the AI2Thor simulator.
-The construction of the knowledge base is discussed in Section 3.3.
-We generated approximately 72K trajectories, with 80% used for training and 20% reserved for validation.
+For the Heuristic Process, we generated a dataset of 72K expert trajectories in AI2THOR.
+An optimal path was computed for each instruction using the A* algorithm (Hart et al., 1968).
+Once the target object was detected in the agent’s field of view, we used a VLM to generate a rationale for the path taken.
+This data was formatted into a VQA structure, where the question (Q) was the target’s location and the answer (A) was the scene description, reasoning, and decision sequence.
+This dataset, used to populate our Knowledge Base, was split into 80% for training and 20% for validation.
 Additional details are provided in Appendix A.3.
 4.2. Implementation Details
-We employ UniMODE (Li et al., 2024) for 3D robot perception and GPT-4 for the Analytic Process to enable rational and logical thinking.
-Qwen2-VL-7B (Wang et al., 2024a) is employed for both the Exploit module in the Heuristic Process and Demand Matching, with multiple LoRA configurations for optimal performance.
-All experiments are trained on 4 RTX6000 GPUs.
-Please refer to Appendix D for further implementation details.
+For the 3D Robot Perception module, we utilize UniMODE (Li et al., 2024).
+The model is trained for 30,000 iterations with a batch size of 64 on four RTX6000 GPUs, a process that takes approximately 10 hours.
+Input images are resized from 448 × 448 to 640 × 512 pixels.
+We use the AdamW optimizer with a learning rate of , and a weight decay of .
+A CosineAnnealing scheduler is used, with a warm-up factor of 0.3333 for the first 2000 iterations.
+To ensure stability, gradient clipping is applied at a threshold of 35.0.
+For the Demand Matching, we conduct SFT on Qwen2-VL-7B (Wang et al., 2024a) with LoRA, training with a batch size of 1 for 5 epochs on 4 RTX6000 GPUs in 1 hour.
+Key hyperparameters include a learning rate of , gradient accumulation of 8, and a cosine scheduler.
+LoRA is applied with a rank of 64 and alpha of 128, while freezing the ViT layers.
+For the Heuristic Process, we conduct SFT on Qwen2-VL-7B for 5 epochs using knowledge base samples, which takes about 14 hours, with consistent hyperparameters from the Demand Matching.
 4.3. Evaluation in Closed-Loop Navigation
 We conduct closed-loop experiments in AI2Thor, a widely-used open-source simulator, to evaluate the performance of CogDDN on 400 scenes using three metrics: Navigation Success Rate (NSR), Navigation Success Rate weighted by the path length (SPL) (Anderson et al., 2018), and Selection Success Rate (SSR).
-Table 2 compares our method with competitive methods offered by DDN (Wang et al., 2023) on the ProcThor benchmark.
+Table 2 compares our method with competitive methods offered by DDN (Wang et al., 2023) on the ProcThor.
 As shown in Table 2, CogDDN outperforms all single-view camera-only methods.
@@
 4.4. Ablation Study
-We conduct ablation studies on Exploit, CoT, and Reflection in a closed-loop navigation system.
+We conduct ablation studies on Exploit CoT, and Reflection in a closed-loop navigation system.
 The experiments were carried out on both unseen scenes and unseen instructions, demonstrating the generalization and continuous learning capabilities of our CogDDN.
@@
 Ablation study of Reflection.
-Reflection plays a crucial role in enhancing the continuous learning capabilities of our proposed CogDDN.
-It involves analyzing mistakes and integrating correct experiences into the knowledge base, thereby promoting self-improvement through proactive summarization and the accumulation of experiences in previously unseen environments.
-In our closed-loop experiment, we conducted four rounds of 500 epochs each, where the experiences generated by the reflection mechanism were added to the knowledge base and used to fine-tune the VLM.
-As shown in Figure 7, after four rounds of iterations, both SSR and NSR showed marginal improvements, as the reflection mechanism allows the agent to navigate around obstacles once they are encountered, resulting in only modest improvements in these metrics.
-However, SPL exhibits a significant increase, as the agent, learning from experiences, anticipates obstacles and takes proactive actions.
-This highlights the continuous learning capability of our CogDDN, where performance improves with experience.
+Reflection is crucial for enhancing the continuous learning capabilities of our proposed CogDDN.
+This mechanism promotes self-improvement by analyzing navigational failures and integrating corrected experiences into the knowledge base, allowing the agent to proactively accumulate knowledge in unseen environments.
+We conducted a closed-loop experiment over four iterative rounds of 500 epochs each to evaluate this capability.
+In each round, experiences generated by the reflection mechanism were used to augment the knowledge base and fine-tune the VLM.
+As shown in Figure 7, after four rounds of iterative learning, both SSR and NSR showed only marginal gains, which is an expected outcome as the reflection mechanism primarily enables the agent to recover from obstacles after they are encountered, leading to only modest improvements in these goal-completion metrics.
+In contrast, the SPL exhibited a significant increase.
+This substantial improvement in SPL demonstrates that as the agent accumulates experience, it learns to anticipate potential obstacles and take proactive detours, resulting in more efficient paths.
+This outcome underscores the effective continuous learning capability of CogDDN, as its performance demonstrably improves with experience.
 Ablation of Reflection.
 5. Conclusion
-In this paper, we present CogDDN, a dual-process closed-loop navigation system that mimics the human attentional mechanism by selectively focusing on key objects necessary for fulfilling user demands.
-We have developed a dual-process decision-making module that simulates human cognitive processes.
-Chain of Thought (CoT) reasoning is incorporated to enhance the decision-making process.
-Moreover, we have integrated a cumulative knowledge base to enable the continuous self-improvement of the algorithm.
-Our evaluations demonstrate that the proposed method achieves remarkable closed-loop performance while requiring limited training data.
+In this paper, we introduced CogDDN, a novel dual-process closed-loop navigation system designed to emulate human attentional mechanisms.
+Central to our approach is a dual-process decision-making module that simulates human cognition, where a deliberative process is enhanced by Chain of Thought (CoT) reasoning to navigate complex scenarios.
+Furthermore, we integrated a cumulative knowledge base that enables the agent to continuously learn and self-improve from its experiences.
+Our comprehensive evaluations demonstrate that CogDDN achieves state-of-the-art performance in closed-loop navigation tasks while requiring minimal training data, underscoring its efficiency and practical potential.
 Limitations and Broader Impacts.
@@
 A.1. Annotations for 3D Robot Perception
-We utilized data collected from the AI2THOR simulator to train the 3D Robot Perception Module.
-The process of obtaining 3D bounding box information involved generating point clouds for each object using segmentation and depth maps provided by the AI2THOR simulator.
-These point clouds were used to extract the minimal bounding box that fully encloses each object, allowing us to derive accurate 3D bounding box information.
+The 3D Robot Perception Module was trained on data from the AI2THOR simulator.
+To derive the 3D bounding box ground truth, we first generated a point cloud for each object using the simulator’s segmentation and depth maps.
+Subsequently, the minimal bounding box enclosing this point cloud was extracted to serve as the final, accurate 3D label.
 Each frame in the dataset contains the following detailed annotations:
@@
 A.2. Annotations for Demand Matching
-In this study, we curated a specialized instruction-attribute dataset to train the Demand Matching Module.
-The dataset construction process involved leveraging the provided instruction data, which was paired with the objects that could potentially fulfill the instructions.
-The process begins with each instruction and its associated objects being input into a Large Language Model (LLM), which generates the common attributes shared by objects that could satisfy the instruction.
-These attributes are crucial for matching objects with the instruction.
-A QA-format dataset is then created, where the first part consists of the common attributes that define the correct object, assisting the Demand Matching Module in identifying the appropriate object.
-The second part includes a list of objects, selected from a predefined set, that align with the generated attributes and are deemed to fulfill the instruction.
-A case of Demand Matching dataset is shown in Figure 8.
+To train the Demand Matching Module, we constructed a specialized instruction-attribute dataset.
+This dataset was built from instruction-object pairs, where each instruction is associated with a set of candidate objects capable of fulfilling its requirements.
+The construction process leverages a Large Language Model (LLM) to extract semantic features.
+For each instruction-object pair, the LLM is tasked with inferring the common attributes shared by the candidate objects that satisfy the instruction.
+These inferred attributes are then formulated into a Question-Answering (QA) format.
+The ”question” component consists of the common attributes defining the target object, while the corresponding ”answer” component contains a list of objects from the candidate set that match these attributes.
+A sample entry from the Demand Matching dataset is illustrated in Figure 8.
 A.3. Annotations for Heuristic Process
-To support the supervised fine-tuning of our Heuristic Process, we created datasets using the AI2THOR simulator.
-For each scenario, we randomly select an instruction and target object, then compute an optimal path using the A* algorithm, which is discretized into a sequence of actions guiding the agent to the target.
-After each action, the simulator produces updated images, used to identify the target object, enabling precise localization.
-When the agent detects the target, VLM generates a structured scene description and reasoning based on the target’s spatial location and navigation decisions.
-As shown in Figure 9, the data is formatted as a Visual Question Answering (VQA) structure, with a question about the information of target object and an answer containing the scene description, reasoning steps, and final decision.
-The structured data is stored in the knowledge base, allowing the system to build a robust understanding of the scene and make informed decisions in future scenarios.
+To generate a dataset for the supervised fine-tuning of our Heuristic Process, we utilized the AI2THOR simulator.
+Each scenario begins with a randomly selected instruction and a corresponding target object.
+We then compute an optimal trajectory to the target using the A* algorithm, which yields a discrete sequence of actions.
+As the agent executes this trajectory, it receives updated visual observations from the simulator at each step, enabling progressive localization of the target.
+Upon successful detection of the object, a Vision-Language Model (VLM) is prompted to generate a structured output.
+This output includes a detailed scene description, a chain of reasoning that rationalizes the navigational choices, and the final decision.
+This information is then formatted into a Visual Question Answering (VQA) structure, as illustrated in Figure 9.
+In this format, a query regarding the target object’s status serves as the ”question,” while the VLM’s generated text constitutes the ”answer.” These structured entries populate a knowledge base, providing the robust data required for our system to learn complex scene understanding and decision-making.
 A case of knowledge base dataset
@@
 It reassesses the previous ’MoveAhead’ decision, suggesting that it may have been based on an inference error, as rotation would have been necessary to avoid colliding with the chair in this situation.
-Appendix D Implementation Details
-We train the UniMODE model with a batch size of 64 for 30,000 iterations on 4 RTX6000 GPUs, taking approximately 10 hours.
-The input image resolution is 448 × 448 pixels, resized to 640 × 512 pixels.
-We use the AdamW optimizer with a learning rate of , weight decay of , and a CosineAnnealing scheduler with a warm-up factor of 0.3333 for the first 2000 iterations.
-Gradient clipping is applied at 35.0 for stability.
-For supervised fine-tuning (SFT), we use the Qwen2-VL-7B model with LoRA, training with a batch size of 1 for five epochs on 4 RTX6000 GPUs in 1 hour.
-Key hyperparameters include a learning rate of , gradient accumulation of 8, and a cosine scheduler.
-LoRA is applied with a rank of 64 and alpha of 128, while freezing the ViT layers.
-For the Heuristic Process, we conduct SFT on Qwen2-VL-7B for five epochs using knowledge base samples, which takes about 14 hours, with consistent hyperparameters from the Demand Matching module.